OpenStudy (anonymous):

Simplify. 4 times the 4th root of 128

OpenStudy (anonymous):

\[4\sqrt[4]{128}\]

OpenStudy (anonymous):

like that right, that is ur question?

OpenStudy (campbell_st):

well 28 = 2^7 \[4 \times \sqrt[4]{2^7} = 4 \times \sqrt[4]{2^4 \times 2^3} = 4 \times \sqrt[4]{2^4} \times \sqrt[4]{2^3}\] this may help..

OpenStudy (campbell_st):

just simplify 2^3 to get the final answer.

OpenStudy (anonymous):

@campbell_st i cant simpfly that

OpenStudy (anonymous):

@some_someone yea thats it

OpenStudy (campbell_st):

2^3 = 2x2x2 you can simplify that

OpenStudy (campbell_st):

this question uses the power of a power index law \[(x^a)^b = x^{a \times b}\] and for radicals \[\sqrt[a]{x} = x^{\frac{1}{a}}\] so this appears to be what you are being asked about.

OpenStudy (campbell_st):

slight error in the 2nd line... if thats a solution...

OpenStudy (campbell_st):

you're missing the radical